# Representing Sugihara monoids

### v3

@@ 3 Constructing finite odd Sugihara chains
    For all $k\in\left\{2,\ldots,n+1-i\right\}$, we have $p_{k}=q_{k}$. Let $r=\left(q_{1},\ldots,q_{n+1-i},q_{n+2-i}-1,\ldots,q_{n}-1\right)$. Then $(r,q)\in L_{n+2-i}$, and so $\left(r^{b},q^{d}\right)\in U_{n+2-i}\subseteq\bigcup_{k=1}^{n+1+j}\subseteq R_{j}$. Note now that we also have $r=\left(p_{1},\ldots,p_{n+1-i},q_{n+2-i}-1,\ldots,q_{n}-1\right)$, so $\left(r^{-b},p^{b}\right)\notin U_{k}$ for all $k\in\left\{1,\ldots,n+1-i\right\}$. Hence, $\left(r^{-b},p^{b}\right)\notin R_{-i}$, which implies that $\left(p^{b},r^{-b}\right)\in\left(R_{-i}\right)^{c\smile}$. It therefore follows that $\left(p^{b},r^{b}\right)\in\left(R_{-i}\right)^{c\smile}\circ\alpha={\sim}R_{-i}=R_{i}$. From this we get $\left(p^{b},q^{d}\right)\in R_{i}\circ R_{j}$.
@@
If $j=-n-1$, then $R_{j}=\varnothing$, and so $R_{i}\circ R_{j}=R_{j}$. If $i=0$, we have $R_{i}=\;\leqslant_{X}$, so $R_{i}\circ R_{j}=R_{j}$ by Lemma 2.4. We now consider the following cases:
Case 1: $-n-1<j<i<0$. Since $R_{i}\subseteq{\leqslant_{X}}$, we get $R_{i}\circ R_{j}\subseteq R_{j}$. Let $\left(p^{b},q^{d}\right)\in R_{j}$. Then $\left(p,q\right)\in L_{k}$ for some $k\in\left\{1,2,\ldots,n+1+j\right\}$. That is, $p_{1}<q_{1}$ or $(p_{1},\dots,p_{k-1})=(q_{1},\dots,q_{k-1})$ and $p_{k}<q_{k}$ for some $k\in\left\{2,\ldots,n+1+j\right\}$. If $p_{1}<q_{1}$, we use the density of $\mathbb{Q}$ to find some $r_{1}\in\mathbb{Q}$ such that $p_{1}<r_{1}<q_{1}$. Then we set $r=\left(r_{1},0\ldots,0\right)$ and obtain $\left(p^{b},r^{b}\right)\in U_{1}\subseteq R_{i}$ and $\left(r^{b},q^{d}\right)\in U_{1}\subseteq R_{j}$. Hence, $\left(p^{b},q^{d}\right)\in R_{i}\circ R_{j}$. Similarly, if $(p_{1},\dots,p_{k-1})=(q_{1},\dots,q_{k-1})$ and $p_{k}<q_{k}$ for some $k\in\left\{2,\ldots,n+1+j\right\}$, we can find some $r_{k}\in\mathbb{Q}$ such that $p_{k}<r_{k}<q_{k}$. Setting $r=\left(q_{1},\ldots,q_{k-1},r_{k},0,\ldots,0\right)$, we obtain $\left(p^{b},r^{b}\right)\in U_{k}\subseteq R_{j}\subseteq R_{i}$ and $\left(r^{b},q^{d}\right)\in U_{k}\subseteq R_{j}$, which means $\left(p^{b},q^{d}\right)\in R_{i}\circ R_{j}$.
Case 2: $i>0,-n-1<j<0$. Since ${\leqslant_{X}}\subseteq R_{i}$, we get $R_{j}\subseteq R_{i}\circ R_{j}$. Let $\left(p^{b},q^{d}\right)\in R_{i}\circ R_{j}$. Then there is some $r^{e}$ such that $\left(p^{b},r^{e}\right)\in R_{i}$ and $\left(r^{e},q^{d}\right)\in R_{j}$. The latter gives $(r,q)\in L_{k}$ for some $k\in\left\{1,\ldots,n+1+j\right\}$. The former gives $\left(p^{b},r^{e}\right)\in{\sim}R_{-i}=\left(R_{-i}\right)^{c\smile}\circ\alpha$, and so $\left(r^{-e},p^{b}\right)\notin R_{-i}$. Hence, $\left(r^{-e},p^{b}\right)\notin U_{\ell}$ for all $\ell\in\left\{1,\ldots,n+1-i\right\}$. That is, $r_{1}\not<p_{1}$ and, for all $\ell\in\left\{2,\ldots,n+1-i\right\}$, if $\left(p_{1},\ldots,p_{\ell-1}\right)=\left(r_{1},\ldots,r_{\ell-1}\right)$ then $r_{\ell}\not<p_{\ell}$. Since $\mathbb{Q}$ is linearly ordered, we get $p_{1}\leqslant r_{1}$ and, for all $\ell\in\left\{2,\ldots,n+1-j\right\}$, if $\left(p_{1},\ldots,p_{\ell-1}\right)=\left(r_{1},\ldots,r_{\ell-1}\right)$ then $p_{\ell}\leqslant r_{1}$. If $r_{1}<q_{1}$, then $p_{1}\leqslant r_{1}<q_{1}$, and so $\left(p^{b},q^{d}\right)\in U_{1}\subseteq R_{j}$. If $p_{1}<r_{1}$, we get $p_{1}<r_{1}\leqslant q_{1}$, so again $\left(p^{b},q^{d}\right)\in U_{1}\subseteq R_{j}$. If $p_{1}=r_{1}=q_{1}$, then $\left(p_{1},\ldots,p_{m-1}\right)=\left(r_{1},\ldots,r_{m-1}\right)=\left(q_{1},\ldots,q_{m-1}\right)$ and $p_{m}<q_{m}$ for some $m\in\left\{2,\ldots,k\right\}$. Hence, $(p,q)\in L_{m}$, and so $\left(p^{b},q^{d}\right)\in U_{m}\subseteq R_{j}$.
@@
    For all $k\in\left\{1,\ldots,n+1-j\right\}$, we have $p_{k}=q_{k}$. Let
    $r=\left(p_{1},\ldots,p_{n+1-j},p_{n+2-j}+1,\ldots,p_{n}+1\right).$
    Then $\left(p,r\right)\in L_{n+2-j}$, and so $\left(p^{b},r^{b}\right)\in U_{n+2-j}\subseteq\bigcup_{k=1}^{n+1+i}U_{k}=R_{i}$. We also have $\left(q,r\right)\notin L_{k}$ for all $k\in\left\{1,\ldots,n+1-j\right\}$. Hence, $\left(q^{-d},r^{b}\right)\notin U_{k}$ for all $k\in\left\{1,\ldots,n+1-j\right\}$, which means $\left(q^{-d},p^{b}\right)\notin R_{-j}$. We thus have $\left(r^{b},q^{-d}\right)\in\left(R_{-j}\right)^{c\smile}$, and thus, $\left(r^{b},q^{d}\right)\in\left(R_{-j}\right)^{c\smile}\circ\alpha={\sim}R_{-j}=R_{j}$. This gives us $\left(p^{b},q^{d}\right)\in R_{i}\circ R_{j}$.
@@
The claim clearly follows for the cases $i=j=0$ and $i=j=-n-1$. So consider the following cases:
Case 1: $-n-1<i=j<0$. Let $\left(p^{b},q^{d}\right)\in R_{i}\circ R_{j}=R_{i}\circ R_{i}$. Then there exists $r^{b}$ such that $\left(p^{b},r^{e}\right)\in R_{i}$ and $\left(r^{e},q^{d}\right)\in R_{i}$. By Lemma 3.10, $R_{i}$ is transitive, so $\left(p^{b},q^{d}\right)\in R_{i}=R_{j}=R_{\text{min}\{i,j\}}$. Now let $\left(p^{b},q^{d}\right)\in R_{\text{min}\{i,j\}}=R_{i}=R_{j}$. Then $(p,q)\in L_{k}$ for some $k\in\left\{1,\ldots,n+1+i\right\}$. There are two cases:
@@
Case 2: $-n-1<i<0$, $j>0$. Since $\leqslant_{X}\;\subseteq R_{j}$, we have $R_{\text{min}\{i,j\}}=R_{i}\subseteq R_{i}\circ R_{j}$. Now let $(p^{b},q^{d})\in R_{i}\circ R_{j}$. Then there is some $r^{b}$ such that $\left(p^{b},r^{e}\right)\in R_{i}$ and $\left(r^{e},q^{d}\right)\in R_{j}$. The former gives $\left(p,r\right)\in L_{k}$ for some $k\in\left\{1,\ldots,n+1+i\right\}$. The latter gives $\left(q^{-d},r^{e}\right)\notin R_{-j}$, and so $(q,r)\notin L_{\ell}$ for all $\ell\in\left\{1,\ldots,n+1-j\right\}$. Hence, $q_{1}\not<r_{1}$ and, for all $\ell\in\left\{2,\ldots,n+1-j\right\}$, if $\left(r_{1},\ldots,r_{\ell-1}\right)=\left(q_{1},\ldots,q_{\ell-1}\right)$ then $q_{\ell}\not<r_{\ell}$. Since $\mathbb{Q}$ is linearly ordered, we obtain $r_{1}\leqslant q_{1}$ and, for all $\ell\in\left\{2,\ldots,n+1-j\right\}$, if $\left(r_{1},\ldots,r_{\ell-1}\right)=\left(q_{1},\ldots,q_{\ell-1}\right)$ then $r_{\ell}\leqslant q_{\ell}$. If $p_{1}<r_{1}$, then $p_{1}<r_{1}\leqslant q_{1}$, so $\left(p^{b},q^{d}\right)\in U_{1}\subseteq R_{i}=R_{\text{min}\{i,j\}}$. If $r_{1}<q_{1}$, we get $p_{1}\leqslant r_{1}<q_{1}$, which implies $\left(p^{b},q^{d}\right)\in U_{1}\subseteq R_{i}=R_{\text{min}\{i,j\}}$. Finally, if $p_{1}=r_{1}=q_{1}$, then $\left(p_{1},\ldots,p_{m-1}\right)=\left(r_{1},\ldots,r_{m-1}\right)=\left(q_{1},\ldots,q_{m-1}\right)$ and $p_{m}<q_{m}$ for some $m\in\left\{2,\ldots,k\right\}$. Hence, $(p,q)\in L_{m}$, and so $\left(p^{b},q^{d}\right)\in U_{m}\subseteq R_{i}=R_{\text{min}\{i,j\}}$.
@@
Case 3: $-n-1<j<0,i>0$. Since $\leqslant_{X}\;\subseteq R_{i}$, we have $R_{\text{min}\{i,j\}}=R_{j}\subseteq R_{i}\circ R_{j}$. Let $(p^{b},q^{d})\in R_{i}\circ R_{j}$. Then there exists some $r^{b}$ such that $\left(p^{b},r^{e}\right)\in R_{i}$ and $\left(r^{e},q^{d}\right)\in R_{j}$. The second part gives $\left(r,q\right)\in L_{\ell}$ for some $\ell\in\left\{1,\ldots,n+1+j\right\}$. The first part gives $\left(r^{-e},p^{b}\right)\notin R_{-i}$, which means $(r,p)\notin L_{k}$ for all $k\in\left\{1,\ldots,n+1-i\right\}$. Hence, $r_{1}\not<p_{1}$ and, for all $k\in\left\{2,\ldots,n+1-i\right\}$, if $\left(r_{1},\ldots,r_{k-1}\right)=\left(p_{1},\ldots,p_{k-1}\right)$ then $r_{k}\not<p_{k}$. Since $\mathbb{Q}$ is linearly ordered, we get $p_{1}\leqslant r_{1}$ and, for all $k\in\left\{2,\ldots,n+1-i\right\}$, if $\left(r_{1},\ldots,r_{k-1}\right)=\left(p_{1},\ldots,p_{k-1}\right)$ then $p_{k}\leqslant r_{k}$. If $r_{1}<q_{1}$, then $p_{1}\leqslant r_{1}<q_{1}$, and so $\left(p^{b},q^{d}\right)\in U_{1}\subseteq R_{j}=R_{\text{min}\{i,j\}}$. If $p_{1}<r_{1}$, then it follows that $p_{1}<r_{1}\leqslant q_{1}$, which gives $\left(p^{b},q^{d}\right)\in U_{1}\subseteq R_{j}=R_{\text{min}\{i,j\}}$. Finally, if $p_{1}=r_{1}=q_{1}$, then $\left(p_{1},\ldots,p_{m-1}\right)=\left(r_{1},\ldots,r_{m-1}\right)=\left(q_{1},\ldots,q_{m-1}\right)$ and $p_{m}<q_{m}$ for some $m\in\left\{2,\ldots,\ell\right\}$. Hence, $(p,q)\in L_{m}$, and so $\left(p^{b},q^{d}\right)\in U_{m}\subseteq R_{j}=R_{\text{min}\{i,j\}}$.
@@
[…]
@@
[…]

### v4

@@ 3 Constructing finite odd Sugihara chains
    For all $k\in\left\{2,\ldots,n+1-i\right\}$, we have $p_{k}=q_{k}$. Let r = (q_1, …, q_n+1-i, q_n+2-i-1,…, q_n-1). Then $(r,q)\in L_{n+2-i}$, and so $\left(r^{b},q^{d}\right)\in U_{n+2-i}\subseteq\bigcup_{k=1}^{n+1+j}\subseteq R_{j}$. Note now that we also have $r=\left(p_{1},\ldots,p_{n+1-i},q_{n+2-i}-1,\ldots,q_{n}-1\right)$, so $\left(r^{-b},p^{b}\right)\notin U_{k}$ for all $k\in\left\{1,\ldots,n+1-i\right\}$. Hence, $\left(r^{-b},p^{b}\right)\notin R_{-i}$, which implies that $\left(p^{b},r^{-b}\right)\in\left(R_{-i}\right)^{c\smile}$. It therefore follows that $\left(p^{b},r^{b}\right)\in\left(R_{-i}\right)^{c\smile}\circ\alpha={\sim}R_{-i}=R_{i}$. From this we get $\left(p^{b},q^{d}\right)\in R_{i}\circ R_{j}$.
@@
If j=n1j=-n-1, then $R_{j}=\varnothing$, and so $R_{i}\circ R_{j}=R_{j}$. If $i=0$, we have $R_{i}=\;\leqslant_{X}$, so $R_{i}\circ R_{j}=R_{j}$ by Lemma 2.4. We now consider the following cases:
Case 1: n1<j<i<0-n-1<j<i<0. Since $R_{i}\subseteq{\leqslant_{X}}$, we get $R_{i}\circ R_{j}\subseteq R_{j}$. Let $\left(p^{b},q^{d}\right)\in R_{j}$. Then $\left(p,q\right)\in L_{k}$ for some $k\in\left\{1,2,\ldots,n+1+j\right\}$. That is, $p_{1}<q_{1}$ or $(p_{1},\dots,p_{k-1})=(q_{1},\dots,q_{k-1})$ and $p_{k}<q_{k}$ for some $k\in\left\{2,\ldots,n+1+j\right\}$. If $p_{1}<q_{1}$, we use the density of $\mathbb{Q}$ to find some $r_{1}\in\mathbb{Q}$ such that $p_{1}<r_{1}<q_{1}$. Then we set $r=\left(r_{1},0\ldots,0\right)$ and obtain $\left(p^{b},r^{b}\right)\in U_{1}\subseteq R_{i}$ and $\left(r^{b},q^{d}\right)\in U_{1}\subseteq R_{j}$. Hence, $\left(p^{b},q^{d}\right)\in R_{i}\circ R_{j}$. Similarly, if $(p_{1},\dots,p_{k-1})=(q_{1},\dots,q_{k-1})$ and $p_{k}<q_{k}$ for some $k\in\left\{2,\ldots,n+1+j\right\}$, we can find some $r_{k}\in\mathbb{Q}$ such that $p_{k}<r_{k}<q_{k}$. Setting $r=\left(q_{1},\ldots,q_{k-1},r_{k},0,\ldots,0\right)$, we obtain $\left(p^{b},r^{b}\right)\in U_{k}\subseteq R_{j}\subseteq R_{i}$ and $\left(r^{b},q^{d}\right)\in U_{k}\subseteq R_{j}$, which means $\left(p^{b},q^{d}\right)\in R_{i}\circ R_{j}$.
Case 2: i>0,n1<j<0i>0,-n-1<j<0. Since ${\leqslant_{X}}\subseteq R_{i}$, we get $R_{j}\subseteq R_{i}\circ R_{j}$. Let $\left(p^{b},q^{d}\right)\in R_{i}\circ R_{j}$. Then there is some $r^{e}$ such that $\left(p^{b},r^{e}\right)\in R_{i}$ and $\left(r^{e},q^{d}\right)\in R_{j}$. The latter gives $(r,q)\in L_{k}$ for some $k\in\left\{1,\ldots,n+1+j\right\}$. The former gives $\left(p^{b},r^{e}\right)\in{\sim}R_{-i}=\left(R_{-i}\right)^{c\smile}\circ\alpha$, and so $\left(r^{-e},p^{b}\right)\notin R_{-i}$. Hence, $\left(r^{-e},p^{b}\right)\notin U_{\ell}$ for all $\ell\in\left\{1,\ldots,n+1-i\right\}$. That is, $r_{1}\not<p_{1}$ and, for all $\ell\in\left\{2,\ldots,n+1-i\right\}$, if $\left(p_{1},\ldots,p_{\ell-1}\right)=\left(r_{1},\ldots,r_{\ell-1}\right)$ then $r_{\ell}\not<p_{\ell}$. Since $\mathbb{Q}$ is linearly ordered, we get $p_{1}\leqslant r_{1}$ and, for all $\ell\in\left\{2,\ldots,n+1-j\right\}$, if $\left(p_{1},\ldots,p_{\ell-1}\right)=\left(r_{1},\ldots,r_{\ell-1}\right)$ then $p_{\ell}\leqslant r_{1}$. If $r_{1}<q_{1}$, then $p_{1}\leqslant r_{1}<q_{1}$, and so $\left(p^{b},q^{d}\right)\in U_{1}\subseteq R_{j}$. If $p_{1}<r_{1}$, we get $p_{1}<r_{1}\leqslant q_{1}$, so again $\left(p^{b},q^{d}\right)\in U_{1}\subseteq R_{j}$. If $p_{1}=r_{1}=q_{1}$, then $\left(p_{1},\ldots,p_{m-1}\right)=\left(r_{1},\ldots,r_{m-1}\right)=\left(q_{1},\ldots,q_{m-1}\right)$ and $p_{m}<q_{m}$ for some $m\in\left\{2,\ldots,k\right\}$. Hence, $(p,q)\in L_{m}$, and so $\left(p^{b},q^{d}\right)\in U_{m}\subseteq R_{j}$.
@@
    For all $k\in\left\{1,\ldots,n+1-j\right\}$, we have $p_{k}=q_{k}$. Let r = (p_1, …, p_n+1-j, p_n+2-j+1, …, p_n + 1). Then $\left(p,r\right)\in L_{n+2-j}$, and so $\left(p^{b},r^{b}\right)\in U_{n+2-j}\subseteq\bigcup_{k=1}^{n+1+i}U_{k}=R_{i}$. We also have $\left(q,r\right)\notin L_{k}$ for all $k\in\left\{1,\ldots,n+1-j\right\}$. Hence, $\left(q^{-d},r^{b}\right)\notin U_{k}$ for all $k\in\left\{1,\ldots,n+1-j\right\}$, which means $\left(q^{-d},p^{b}\right)\notin R_{-j}$. We thus have $\left(r^{b},q^{-d}\right)\in\left(R_{-j}\right)^{c\smile}$, and thus, $\left(r^{b},q^{d}\right)\in\left(R_{-j}\right)^{c\smile}\circ\alpha={\sim}R_{-j}=R_{j}$. This gives us $\left(p^{b},q^{d}\right)\in R_{i}\circ R_{j}$.
@@
The claim clearly follows for the cases $i=j=0$ and i=j=n1i=j=-n-1. So consider the following cases:
Case 1: n1<i=j<0-n-1<i=j<0. Let $\left(p^{b},q^{d}\right)\in R_{i}\circ R_{j}=R_{i}\circ R_{i}$. Then there exists $r^{b}$ such that $\left(p^{b},r^{e}\right)\in R_{i}$ and $\left(r^{e},q^{d}\right)\in R_{i}$. By Lemma 3.10, $R_{i}$ is transitive, so $\left(p^{b},q^{d}\right)\in R_{i}=R_{j}=R_{\text{min}\{i,j\}}$. Now let $\left(p^{b},q^{d}\right)\in R_{\text{min}\{i,j\}}=R_{i}=R_{j}$. Then $(p,q)\in L_{k}$ for some $k\in\left\{1,\ldots,n+1+i\right\}$. There are two cases:
@@
Case 2: n1<i<0-n-1<i<0, $j>0$. Since $\leqslant_{X}\;\subseteq R_{j}$, we have $R_{\text{min}\{i,j\}}=R_{i}\subseteq R_{i}\circ R_{j}$. Now let $(p^{b},q^{d})\in R_{i}\circ R_{j}$. Then there is some $r^{b}$ such that $\left(p^{b},r^{e}\right)\in R_{i}$ and $\left(r^{e},q^{d}\right)\in R_{j}$. The former gives $\left(p,r\right)\in L_{k}$ for some $k\in\left\{1,\ldots,n+1+i\right\}$. The latter gives $\left(q^{-d},r^{e}\right)\notin R_{-j}$, and so $(q,r)\notin L_{\ell}$ for all $\ell\in\left\{1,\ldots,n+1-j\right\}$. Hence, $q_{1}\not<r_{1}$ and, for all $\ell\in\left\{2,\ldots,n+1-j\right\}$, if $\left(r_{1},\ldots,r_{\ell-1}\right)=\left(q_{1},\ldots,q_{\ell-1}\right)$ then $q_{\ell}\not<r_{\ell}$. Since $\mathbb{Q}$ is linearly ordered, we obtain $r_{1}\leqslant q_{1}$ and, for all $\ell\in\left\{2,\ldots,n+1-j\right\}$, if $\left(r_{1},\ldots,r_{\ell-1}\right)=\left(q_{1},\ldots,q_{\ell-1}\right)$ then $r_{\ell}\leqslant q_{\ell}$. If $p_{1}<r_{1}$, then $p_{1}<r_{1}\leqslant q_{1}$, so $\left(p^{b},q^{d}\right)\in U_{1}\subseteq R_{i}=R_{\text{min}\{i,j\}}$. If $r_{1}<q_{1}$, we get $p_{1}\leqslant r_{1}<q_{1}$, which implies $\left(p^{b},q^{d}\right)\in U_{1}\subseteq R_{i}=R_{\text{min}\{i,j\}}$. Finally, if $p_{1}=r_{1}=q_{1}$, then $\left(p_{1},\ldots,p_{m-1}\right)=\left(r_{1},\ldots,r_{m-1}\right)=\left(q_{1},\ldots,q_{m-1}\right)$ and $p_{m}<q_{m}$ for some $m\in\left\{2,\ldots,k\right\}$. Hence, $(p,q)\in L_{m}$, and so $\left(p^{b},q^{d}\right)\in U_{m}\subseteq R_{i}=R_{\text{min}\{i,j\}}$.
@@
Case 3: n1<j<0,i>0-n-1<j<0,i>0. Since $\leqslant_{X}\;\subseteq R_{i}$, we have $R_{\text{min}\{i,j\}}=R_{j}\subseteq R_{i}\circ R_{j}$. Let $(p^{b},q^{d})\in R_{i}\circ R_{j}$. Then there exists some $r^{b}$ such that $\left(p^{b},r^{e}\right)\in R_{i}$ and $\left(r^{e},q^{d}\right)\in R_{j}$. The second part gives $\left(r,q\right)\in L_{\ell}$ for some $\ell\in\left\{1,\ldots,n+1+j\right\}$. The first part gives $\left(r^{-e},p^{b}\right)\notin R_{-i}$, which means $(r,p)\notin L_{k}$ for all $k\in\left\{1,\ldots,n+1-i\right\}$. Hence, $r_{1}\not<p_{1}$ and, for all $k\in\left\{2,\ldots,n+1-i\right\}$, if $\left(r_{1},\ldots,r_{k-1}\right)=\left(p_{1},\ldots,p_{k-1}\right)$ then $r_{k}\not<p_{k}$. Since $\mathbb{Q}$ is linearly ordered, we get $p_{1}\leqslant r_{1}$ and, for all $k\in\left\{2,\ldots,n+1-i\right\}$, if $\left(r_{1},\ldots,r_{k-1}\right)=\left(p_{1},\ldots,p_{k-1}\right)$ then $p_{k}\leqslant r_{k}$. If $r_{1}<q_{1}$, then $p_{1}\leqslant r_{1}<q_{1}$, and so $\left(p^{b},q^{d}\right)\in U_{1}\subseteq R_{j}=R_{\text{min}\{i,j\}}$. If $p_{1}<r_{1}$, then it follows that $p_{1}<r_{1}\leqslant q_{1}$, which gives $\left(p^{b},q^{d}\right)\in U_{1}\subseteq R_{j}=R_{\text{min}\{i,j\}}$. Finally, if $p_{1}=r_{1}=q_{1}$, then $\left(p_{1},\ldots,p_{m-1}\right)=\left(r_{1},\ldots,r_{m-1}\right)=\left(q_{1},\ldots,q_{m-1}\right)$ and $p_{m}<q_{m}$ for some $m\in\left\{2,\ldots,\ell\right\}$. Hence, $(p,q)\in L_{m}$, and so $\left(p^{b},q^{d}\right)\in U_{m}\subseteq R_{j}=R_{\text{min}\{i,j\}}$.
@@
where $\Rightarrow$ is defined, as in the abstract case, in terms of $\cap$, $\cup$ and $\sim$}.
@@
The fact that the ordered set forms a chain is Lemma 3.8. Comparing Proposition 3.18 with the definition of $\mathbf{S}_{2n+3}$ in Section 2.1 shows that the monoid operation is defined as it should be for an odd Sugihara chain. The fact that $\sim$} is order-reversing follows from the fact that $R_{i}={\sim}R_{-i}$ and Lemma 3.8. The isomorphism is given by $\varphi:S_{2n+3}\to\mathcal{R}_{2n+3}$ where $\varphi(a_{i})=R_{i}$.